# From South to North: Leveraging Ground-Based LATs for Full-Sky CMB Delensing and Constraints on $r$

### v1

@@ Abstract
Delensing—the process of mitigating the lensing-induced B-mode contamination in cosmic microwave background (CMB) observations—will be a pivotal challenge for next-generation CMB experiments seeking to detect primordial gravitational waves (PGWs) through B-mode polarization. This process requires an accurate lensing tracer, which can be obtained either through internal reconstruction from high-resolution CMB observations or from external large-scale structure (LSS) surveys. Ground-based large-aperture telescopes (LATs) are crucial for internal reconstruction, yet existing and planned facilities are confined to the southern hemisphere, limiting effective delensing to that region. In this work, we assess the impact of introducing a northern hemisphere LAT, assumed to be situated near AliCPT (hence termed Ali-like LAT, or LATN), on delensing performance and PGW detection, using simulations. Our baseline setup includes a space-based small-aperture mission (LiteBIRD-like, SAT) and a southern LAT (SO-like, LATS). External LSS tracers, which have been shown to play an important role in delensing before the availability of ultra-sensitive polarization data, are also considered. We find that southern-hemisphere internal delensing reduces the uncertainty in $r$ by $\sim$21compared to the no-delensing case. Adding LATN enables full-sky internal delensing, achieving a further $\sim$19reduction—comparable to that from including LSS tracers ($\sim$17%). Once LATN is included, the marginal benefit of LSS tracers drops to $\sim$10%. These results highlight the significant role of LATN in advancing delensing capabilities and improving PGW constraints.
@@ 1 Introduction
The removal of lensing-induced contamination from CMB anisotropy maps has been a major focus in recent literature [23, 24, 25, 26]. Accurate delensing requires prior knowledge of the lensing potential $\phi$, which can be obtained through various methods in practice. One approach involves internal reconstruction using quadratic estimators based on CMB two-point statistics [27], while Bayesian iterative methods [28, 29] provide a statistically optimal alternative. Additionally, external mass tracers, such as the Cosmic Infrared Background and high-redshift galaxy surveys, act as effective proxies for lensing and have shown significant improvements, particularly in the absence of future polarization surveys with deeper sensitivities [30, 31, 32]. For the internal lensing reconstruction, high-resolution ($-\ell$) observation with low instrumental noise is necessary to reduce reconstruction uncertainty. Ground-based telescopes are exceptionally well-suited for this task, as their large apertures—readily achievable in ground-based installations—enable precise measurements of small-scale CMB fluctuations. Although atmospheric noise typically degrades large-scale sensitivity, its impact on lensing reconstruction is limited if properly accounted for. Furthermore, foreground contamination across multiple frequency channels (27–280,GHz) can be effectively mitigated using the Needlet Internal Linear Combination (NILC) method.
@@
For future ultra-low-noise observations [29], internal lensing reconstruction is expected to provide more accurate lensing estimates. In this context, achieving a full-sky internal reconstruction of the lensing potential using ground-based large-aperture telescopes (LATs) is crucial for enabling full-sky delensing and tightening constraints on the tensor-to-scalar ratio $r$. In this work, to emphasize the importance of full-sky internal delensing, we focus on a near-future achievable configuration, using a southern LAT (SO-like, denoted as LATS) as the baseline, and assess the additional contribution of a northern LAT, assumed to be situated near AliCPT[35] (hence termed Ali-like LAT, or LATN), to full-sky primordial gravitational wave (PGW) detection through delensing. Our simulations combine polarization data from a hypothetical satellite mission (SAT)—featuring ultra-low instrumental noise, no atmospheric contamination, and wide frequency coverage (40–402 GHz across 15 channels)—with data from both LATs. Foreground removal is performed using the NILC method. We demonstrate that, even with a near-future LAT configurations, the inclusion of LATN, enabling full-sky internal delensing, yields a substantial improvement in constraints on $r$—comparable to that achieved through the use of external large-scale structure tracers.
This paper is organized as follows. In Section 2, we introduce the delensing methodology along with the associated debiasing formalism. Section 3 outlines the complete analysis pipeline, including map simulation, foreground cleaning, map co-addition, lensing reconstruction, CMB B-mode delensing, and parameter estimation. In Section 4, we present the performance of each step in the pipeline, culminating in the final parameter constraints. We conclude the paper in Section 5.
@@ 2 Large-scale B-mode Delensing
$\tilde{\Theta}(\hat{\mathbf{n}})$ $=\Theta(\hat{\mathbf{n}}+\mathbf{d}(\hat{\mathbf{n}}))=\Theta(\hat{\mathbf{n}})+\nabla_{i}\Theta(\hat{\mathbf{n}})\mathbf{d}^{i}(\hat{\mathbf{n}})+\mathcal{O}(\phi^{2}),$ (2.1)
$[\tilde{Q}\pm i\tilde{U}](\hat{\mathbf{n}})$ $=[Q\pm iU](\hat{\mathbf{n}}+\mathbf{d}(\hat{\mathbf{n}}))=[Q\pm iU](\hat{\mathbf{n}})+\nabla_{i}[Q\pm iU](\hat{\mathbf{n}})\mathbf{d}^{i}(\hat{\mathbf{n}})+\mathcal{O}(\phi^{2}),$ (2.2)
where tildes indicate lensed quantities and $\mathbf{d}$ is the deflection angle, given by the gradient of the lensing potential $\mathbf{d}(\hat{\mathbf{n}})=\nabla\phi(\hat{\mathbf{n}})$ (here we ignore the curl modes) [37], where the lensing potential is the projection of gravitational potential along line-of-sight [27]:
@@
where $p^{+}(p^{-})$ is unity if $\ell+L+\ell^{\prime}$ is even (odd) and zero otherwise. Lensing induced mode coupling for $s$ fields can be expressed as:
@@
×$\times\sqrt{\frac{(2l+1)(2l^{\prime}+1)(2L+1)}{16\pi}}\begin{pmatrix}\ell&\ell^{\prime}&0\\ -s&s^{\prime}&0\end{pmatrix}$
@@ 2.1 Delensing Method Description
where the covariant derivative $\nabla_{i}$ operating on the $s$ field piece of a tensor is equivalent to a gradient operation $D_{i}$ on its $s$ weighted representation, $\sharp$ and $\flat$ denote the ladder operators, and the complex-conjugated vectors $\bar{m}$ and $m$ serve as the basis.
@@
where the primary CMB fields can be recovered from the lensed field by remapping the latter using the inverse deflection angle $\mathbf{d^{inv}}(\hat{\mathbf{n}})$:
$\tilde{X}(\hat{\mathbf{n}})=X(\hat{\mathbf{n}}+\mathbf{d}(\hat{\mathbf{n}}))\Leftrightarrow X(\hat{\mathbf{n}})=\tilde{X}(\hat{\mathbf{n}}+\mathbf{d^{inv}}(\hat{\mathbf{n}})).$ (2.8)
In practice, at low-order approximation, the inverse deflection angle is approximately the negative of the deflection angle. A more accurate estimate can be obtained by solving Eq. (2.7) using a Newton-Raphson scheme, where the inverse deflection angle $\mathbf{d^{\text{inv}}}(\hat{\mathbf{n}})$ is iteratively computed, as described in [25]. The lensing B-mode template is then constructed by subtracting the inverse-lensed B-mode map from the observed data.
@@ 2.2.1 Bias Analysis of the Gradient-Order Template Method
=$&\mathcal{B}^{(1)}[\mathcal{W}^{E}E^{\text{lens}}\ast\mathcal{W}^{\phi}\phi]+\left\{\mathcal{B}^{(1)}[\mathcal{W}^{E}E^{\text{res,NILC}}\ast\mathcal{W}^{\phi}\phi]+\mathcal{B}^{(1)}[\mathcal{W}^{E}(E^{\text{lens}}+E^{\text{res,NILC}})\ast\mathcal{W}^{\phi}\phi^{\text{noise}}]\right\}\end{aligned}$
@@
where the first term represents the signal component of the template, while the remaining terms correspond to noise, foreground residuals, and lensing reconstruction noise. Notably, all of these contributions are suppressed by the Wiener filter.
The auto-power spectrum of the lensing template from Eq. (2.10) is then:
$C^{\text{temp}}=C_{S}^{\text{temp}}+C_{N}^{\text{temp}}+2\langle B^{\text{temp}}_{S}B^{\text{temp}}_{N}\rangle.$ (2.11)
The cross-power spectrum between $B^{\text{temp}}=\mathcal{T}^{-1}B^{\text{lens}}+B_{N}^{\text{temp}}$ and $B^{\text{obs}}=B^{\text{lens}}+B^{\text{res,NILC}}$ is:
$C^{\text{cross}}=\mathcal{T}^{-1}C^{\text{lens}}+\langle B^{\text{lens}}B_{N}^{\text{temp}}\rangle+\langle B^{\text{res,NILC}}B^{\text{temp}}\rangle,$ (2.12)
where $\mathcal{T}$ is the transfer function that compensates for signal attenuation due to the Wiener filter.
In practice, we find that most of the bias terms are realization-independent. Therefore, to estimate these significant bias terms, we can either compute them directly through simulations or extract them by taking the difference between the baseline result and the results obtained using the signal-only simulations.
@@ 2.2.2 Bias Analysis of the Inverse-Lensing Method
$\approx\left\{\mathcal{B}[\mathcal{W}^{E}B^{\text{lens}}\star\mathbf{d^{inv}_{\phi}}]\right\}+\left\{\mathcal{B}[\mathcal{W}^{E}B^{\text{lens}}\star\mathbf{d^{inv}_{\text{noise}}}]+\mathcal{B}[\mathcal{W}^{E}B^{\text{res,NILC}}\star\mathbf{d^{inv}}]-\mathcal{W}^{E}B^{\text{lens}}\right\}$
@@
The lensing template is obtained by subtracting the delensed B-mode from the filtered observed B-modes:
@@
$=B^{\text{temp}}_{S}+B^{\text{df}},$
where the same filter is applied to the observed B-modes to ensure consistency with the delensed B-modes, which experience signal attenuation due to the Wiener filter.
Next, we calculate the auto-power spectrum of the lensing template, given by $B^{\text{temp}}=\mathcal{T}^{-1}B^{\text{lens}}+B^{\text{df}}$:
$C^{\text{temp}}$ $=\mathcal{T}^{-2}C^{\text{lens}}+\langle B^{\text{df}}B^{\text{df}}\rangle+2\mathcal{T}^{-1}\langle B^{\text{lens}}B^{\text{df}}\rangle,$ (2.15)
where $\mathcal{T}$ represents the transfer function used to compensate for signal attenuation caused by the Wiener filter.
@@
$C^{\text{cross}}$ $=\mathcal{T}^{-1}C^{\text{lens}}+\langle B^{\text{lens}}B^{\text{df}}\rangle+\langle B^{\text{res,NILC}}B^{\text{temp}}\rangle.$ (2.16)
These bias terms essentially correspond to those discussed in the previous section, with the exception of some minor high-order effects that depend on the accuracy of the inverse deflection angle $\mathbf{d^{inv}}(\hat{\mathbf{n}})$. Therefore, they can be estimated in a similar manner to the latter.
@@ 3.1 Data Simulation
We employ simulated data to evaluate the feasibility of the delensing pipeline. Specifically, we simulate future CMB observations from various experiments, including a large-aperture, ground-based 666-m telescope in the Northern Hemisphere (LATN), a large-aperture, ground-based 666-m telescope in the Southern Hemisphere (LATS), and a small-aperture 111-m satellite telescope (SAT). We then assess the effectiveness of different delensing methods in reducing the uncertainty in $r$ as shown in Section 4.3.
@@
Figure 2: Survey windows for each observation in Galactic coordinates used in the analysis. For the coadded case, we combine two ground-based LATs with the satellite SAT. The yellow region represents the overlap between SAT and one of the LATs, while the brown region indicates the common region covered by all three observations.
@@
The input unlensed CMB maps are Gaussian realizations generated from a specific power spectrum obtained using the Boltzmann code CAMB [39], based on the Planck 2018 best-fit cosmological parameters [40] with a tensor-to-scalar ratio $r=0$. The Gaussian realization of the lensing potential maps is also generated from the fiducial lensing potential power spectrum. We then utilize the Lenspyx [41, 42] package, which implements an algorithm to distort the primordial signal based on the lensing potential map generated above.
@@
We include both Galactic and extragalactic foregrounds in our simulation using PySM3 [43, 44]. PySM3 generates full-sky simulations of Galactic emissions in intensity and polarization based on realistic space observations (e.g., Planck, WMAP). The modeled diffuse Galactic emissions include thermal dust (d9 model), synchrotron (s4 model), spinning dust (a1 model), and free-free emission (f1 model). The extragalactic emissions incorporated in our simulations include the Sunyaev-Zel’dovich (SZ) effect (ksz1 and tsz1 models) and the cosmic infrared background (cib1 model).
Since the foreground templates described above are derived from real observational data, only a single simulation is available. Consequently, our analysis does not account for uncertainties in foreground modeling. However, we apply the NILC method to clean the observations (see Section 3.2), and the foreground residuals are expected to be largely suppressed. Therefore, we anticipate that our results are not significantly overestimated.
@@
LAT Temperature LAT Polarization Freq. [GHz] FWHM [arcmin] $\sigma_{\text{white}}$ [$\mu$K-arcmin] $N_{\text{red}}\cdot t$ [$\mu$K2-s] $\ell_{\text{knee}}$ $\alpha_{\text{knee}}$ $N_{\text{red}}$ [$\mu$K2] $\ell_{\text{knee}}$ $\alpha_{\text{knee}}$ 27 7.4 52 100 1000 -3.5 $N_{\text{white,pol.}}$ 700 -1.4 39 5.1 27 39 1000 -3.5 $N_{\text{white,pol.}}$ 700 -1.4 93 2.2 5.8 230 1000 -3.5 $N_{\text{white,pol.}}$ 700 -1.4 145 1.4 6.3 1500 1000 -3.5 $N_{\text{white,pol.}}$ 700 -1.4 225 1.0 15 17000 1000 -3.5 $N_{\text{white,pol.}}$ 700 -1.4 280 0.9 37 31000 1000 -3.5 $N_{\text{white,pol.}}$ 700 -1.4
@@
Freq. [GHz] FWHM [arcmin]
Noise level [$\mu$K-arcmin]

### v2

@@ Abstract
Delensing—the process of mitigating the lensing-induced B-mode contamination in cosmic microwave background (CMB) observations—will be a pivotal challenge for next-generation CMB experiments seeking to detect primordial gravitational waves (PGWs) through B-mode polarization. This process requires an accurate lensing tracer, which can be obtained either through internal reconstruction from high-resolution CMB observations or from external large-scale structure (LSS) surveys. Ground-based large-aperture telescopes (LATs) are crucial for internal reconstruction, yet existing and planned facilities are confined to the southern hemisphere, limiting effective delensing to that region. In this work, we assess the impact of introducing a northern hemisphere LAT, assumed to be situated near AliCPT (hence termed Ali-like LAT, or LATN), on delensing performance and PGW detection, using simulations. Our baseline setup includes a space-based small-aperture mission (LiteBIRD-like, SAT) and a southern LAT (SO-like, LATS). External LSS tracers, which have been shown to play an important role in delensing before the availability of ultra-sensitive polarization data, are also considered. We find that southern-hemisphere internal delensing reduces the uncertainty in $r$ by $\sim$17% compared to the no-delensing case. Adding LATN enables full-sky internal delensing, achieving a further $\sim$18% reduction—comparable to that from including LSS tracers ($\sim$13%). Once LATN is included, the marginal benefit of LSS tracers drops to $\sim$10%. These results highlight the significant role of LATN in advancing delensing capabilities and improving PGW constraints.
@@ 1 Introduction
The removal of lensing-induced contamination from CMB anisotropy maps has been a major focus in recent literature [39, 40, 26, 15]. Accurate delensing requires prior knowledge of the lensing potential $\phi$, which can be obtained through various methods in practice. One approach involves internal reconstruction using quadratic estimators based on CMB two-point statistics [86], while Bayesian iterative methods [27, 21] provide a statistically optimal alternative. Additionally, external mass tracers, such as the Cosmic Infrared Background and high-redshift galaxy surveys, act as effective proxies for lensing and have shown significant improvements, particularly in the absence of future polarization surveys with deeper sensitivities [106, 104, 105]. For the internal lensing reconstruction, high-resolution (high-$\ell$) observation with low instrumental noise is necessary to reduce reconstruction uncertainty. Ground-based telescopes are exceptionally well-suited for this task, as their large apertures—readily achievable in ground-based installations—enable precise measurements of small-scale CMB fluctuations. Although atmospheric noise typically degrades large-scale sensitivity, its impact on lensing reconstruction is limited if properly accounted for. Furthermore, foreground contamination across multiple frequency channels (27–280,GHz) can be effectively mitigated using the Needlet Internal Linear Combination (NILC) method.
@@
For future ultra-low-noise observations [21], internal lensing reconstruction is expected to provide more accurate lensing estimates. In this context, achieving a full-sky internal reconstruction of the lensing potential using ground-based large-aperture telescopes (LATs) is crucial for enabling full-sky delensing and tightening constraints on the tensor-to-scalar ratio $r$. In this work, to emphasize the importance of full-sky internal delensing, we focus on a near-future achievable configuration, using a southern LAT (SO-like, denoted as LATS) as the baseline, and assess the additional contribution of a northern LAT, assumed to be situated near AliCPT[69] (hence termed Ali-like LAT, or LATN), to full-sky primordial gravitational wave (PGW) detection through delensing. Our simulations combine polarization data from a hypothetical satellite mission (SAT)—featuring ultra-low instrumental noise, no atmospheric contamination, and wide frequency coverage (40–402 GHz across 15 channels)—with data from both LATs. Foreground removal is performed using the NILC method. We demonstrate that, even with a near-future LAT configurations, the inclusion of LATN, enabling full-sky internal delensing, yields a substantial improvement in constraints on rr—comparable to that achieved through the use of external large-scale structure tracers.
This paper is organized as follows. In Section 2, we introduce the delensing methodology along with the associated debiasing formalism. Section 3 outlines the complete analysis pipeline, including map simulation, foreground cleaning, map co-addition, lensing reconstruction, CMB B-mode delensing, and parameter estimation. In Section 4, we present the performance of each step in the pipeline, culminating in the final parameter constraints. We conclude the paper in Section 5. Further discussion and clarification of the delensing methods are provided in Appendix A. Additional technical details relevant to the likelihood analysis are presented in Appendix B. The contribution from the LT at locations without LAT observations is discussed in Appendix C. An alternative choice of lensing estimators for mitigating extragalactic foregrounds is presented in Appendix D.
@@ 2 Large-scale B-mode Delensing
where $p^{+}(p^{-})$ is unity if $\ell+L+\ell^{\prime}$ is even (odd) and zero otherwise. Lensing induced mode coupling for spin-$s$ fields can be expressed as:
@@
$\times\sqrt{\frac{(2l+1)(2l^{\prime}+1)(2L+1)}{16\pi}}\begin{pmatrix}\ell&\ell^{\prime}&0\\ -s&s^{\prime}&0\end{pmatrix}$.
@@ 2.1 Delensing Method Description
where the covariant derivative $\nabla_{i}$ operating on the spin-$s$ field piece of a tensor is equivalent to a gradient operation $D_{i}$ on its spin-$s$ weighted representation, $\sharp$ and $\flat$ denote the ladder operators, and the complex-conjugated vectors $\bar{m}$ and $m$ serve as the basis.
@@
[…]
$).\tilde{X}(\hat{\mathbf{n}})=X(\hat{\mathbf{n}}+\mathbf{d}(\hat{\mathbf{n}}))\Leftrightarrow X(\hat{\mathbf{n}})=\tilde{X}(\hat{\mathbf{n}}+\mathbf{d^{inv}}(\hat{\mathbf{n}}$)). (2.8)
[…]
@@ 2.2.1 Bias Analysis of the Gradient-Order Template Method
$=\begin{aligned} &\mathcal{B}^{(1)}[\mathcal{W}^{E}E^{\text{lens}}\ast\mathcal{W}^{\phi}\phi]+\left\{\mathcal{B}^{(1)}[\mathcal{W}^{E}E^{\text{res,NILC}}\ast\mathcal{W}^{\phi}(\phi+\phi^{\text{noise}})]+\mathcal{B}^{(1)}[\mathcal{W}^{E}E^{\text{lens}}\ast\mathcal{W}^{\phi}\phi^{\text{noise}}]\right\}\end{aligned}$
@@
where $\mathcal{B}^{(1)}[E\ast\phi]$ represents the operation of constructing the gradient-order template using E-mode field and $\phi$. Here, the first term represents the signal component of the template, while the noise arises from two distinct sources: (i) foreground residuals after NILC processing, which contaminate the lensing template through the reconstruction pipeline; and (ii) the spurious lensing effect on the true $E$-modes induced by the lensing reconstruction noise.
@@ 2.2.2 Bias Analysis of the Inverse-Lensing Method
$\approx\left\{\mathcal{B}[\mathcal{W}^{E}B^{\text{lens}}\star\mathbf{d^{inv}_{\phi}}]\right\}+\left\{[\mathcal{B}[\mathcal{W}^{E}B^{\text{lens}}\star\mathbf{d^{inv}_{\text{noise}}}]-\mathcal{W}^{E}B^{\text{lens}}]+\mathcal{B}[\mathcal{W}^{E}B^{\text{res,NILC}}\star\mathbf{d^{inv}}]+B^{\text{de,hi}}\right\}$
@@
where $\mathcal{B}[B\star\mathbf{d^{inv}}]$ represents the inverse remapping operation of lensed B mode with $\mathbf{d^{inv}}$ to obtain the delensed B-modes. Notice that we have applied a linear approximation when moving from the second row to the third, and all higher-order terms neglected in this step are absorbed into $B^{\mathrm{de,hi}}$. This should be distinguished from the higher-order lensing contributions contained in the term $\mathcal{B}[\,B\star\mathbf{d}^{\mathrm{inv}}]$. The lensing B-mode template:
@@
$=[\mathcal{W}^{E}B^{\text{lens}}-B^{\text{del}}_{S}]+\left\{\begin{aligned} &\mathcal{W}^{E}B^{\text{res,NILC}}-\mathcal{B}(\mathcal{W}^{E}B^{\text{res,NILC}}\star\mathbf{d^{inv}})\\ &\quad+[\mathcal{W}^{E}B^{\text{lens}}-\mathcal{B}(\mathcal{W}^{E}B^{\text{lens}}\star\mathbf{d^{inv}_{\text{noise}}})]-B^{\text{de,hi}}\end{aligned}\right\}$
$=B^{\text{temp}}_{S}+B_{N}^{\text{temp}},$
where $B_{N}^{\mathrm{temp}}$ corresponds to $B^{\mathrm{temp}}_{N}$ in Eq. 2.10. For both the Gradient-Order Template Method and the Inverse-Lensing Method, we find that the noise terms arise from similar physical origins, although the latter includes additional higher-order contributions through $B^{\mathrm{de,hi}}$ beyond the gradient approximation. A comparison of these terms for the two methods is shown in Figure 13. Overall, the two approaches exhibit consistent behavior on large angular scales.
2.2.3 Bias Analysis Formula on power spectral level
With the unified notation for the terms in Eq. 2.10 and Eq. 2.12, we proceed to compute the power spectra of the LT and the observed maps. These spectra form the data vector used in the likelihood analysis to obtain constraints on $r$. The auto-power spectrum of the lensing template, given by $B^{\text{temp}}=B^{\text{temp}}_{S}+B_{N}^{\text{temp}}$, is:
$C^{\text{temp}}$ $=\langle B^{\text{temp}}_{S}B^{\text{temp}}_{S}\rangle+\langle B_{N}^{\text{temp}}B_{N}^{\text{temp}}\rangle+2\langle B^{\text{temp}}_{S}B_{N}^{\text{temp}}\rangle.$ (2.13)
@@ 2.2.3 Bias Analysis Formula on power spectral level
$C^{\text{cross}}$ $=\langle B^{\text{lens}}B_{S}^{\text{temp}}\rangle+\langle B^{\text{lens}}B_{N}^{\text{temp}}\rangle+\langle B^{\text{res,NILC}}B^{\text{temp}}\rangle.$ (2.14)
We present a detailed discussion of the comparison and origin of these terms in Appendix A. The cross-bias terms discussed above are expected to be consistent with zero when either large-scale polarization modes are removed from the internal lensing reconstruction or an external lensing proxy is used. In contrast, $\langle B_{N}^{\mathrm{temp}}B_{N}^{\mathrm{temp}}\rangle$ is found to be of the same order as the CMB lensing $B$-mode signal. Hence, this auto-power term represents the dominant bias that must be carefully accounted for to avoid introducing significant bias in the likelihood analysis.
@@ 3.1 Data Simulation
We employ simulated data to evaluate the feasibility of the delensing pipeline. Specifically, we simulate future CMB observations from various experiments, including a large-aperture, ground-based $6$-m telescope in the Northern Hemisphere (LATN), a large-aperture, ground-based $6$-m telescope in the Southern Hemisphere (LATS), and a small-aperture $1$-m satellite telescope (SAT). We then assess the effectiveness of different delensing methods in reducing the uncertainty in $r$ as shown in Section 4.3.
@@
Figure 2: Survey windows for each observation in Celestial coordinates used in the analysis. For the coadded case, we combine two ground-based LATs with the satellite SAT. The yellow region represents the overlap between SAT and one of the LATs, while the brown region indicates the common region covered by all three observations.
@@
We include both Galactic and extragalactic foregrounds in our simulations. For the Galactic foregrounds, we use PySM3 [125, 113], which generates full-sky simulations of Galactic emissions in both intensity and polarization based on realistic observational data (e.g., Planck and WMAP). The modeled diffuse Galactic emissions include thermal dust (d9 model), synchrotron (s4 model), spinning dust (a1 model), and free-free emission (f1 model). Since these foreground templates are derived from actual observations, only a single realization is available for each component.
For extragalactic foregrounds—including tSZ, kSZ, the CIB, and radio point sources—these components are intrinsically correlated with the lensing potential because they trace the same underlying matter distribution. In addition, they exhibit strong non-Gaussianity on small angular scales, primarily due to the discrete and clustered nature of the underlying astrophysical sources, with additional contributions from the non-linear evolution of large-scale structure. Both the correlation and the non-Gaussianity introduce biases and increase the variance in CMB lensing reconstruction, and numerous works have proposed strategies to mitigate these effects [89, 96, 98, 97, 73, 71]. These foreground-induced systematics propagate into the reconstructed lensing map and, consequently, into the delensing procedure, ultimately leading to increased bias and uncertainty in the inferred constraint on $r$ [17]. Several methods have been proposed to mitigate extragalactic contamination in CMB lensing reconstruction. The standard internal linear combination (ILC) method can leave point-like residuals, such as from point sources or tSZ, but known foregrounds can be deprojected at the cost of increased map variance [64]. Modified lensing estimators, including bias-hardened estimators [84, 89, 97], shear estimators [90, 98] and symmetric multi-frequency–cleaned estimators [38, 74, 37], further reduce foreground biases, albeit with some loss in signal-to-noise. Since techniques for mitigating biases from extragalactic foregrounds in lensing reconstruction are already well established, we do not address this issue in detail in the present work. An alternative foreground-mitigation approach is briefly discussed in Appendix D.
To fully capture the influence of extragalactic foregrounds, multiple realizations are required in the analysis pipeline. Realistic correlated mock data are typically generated using N-body simulations, which model dark matter halos through the gravitational interactions of a large number of particles. These halos form the scaffolding for galaxies and other cosmic structures. Major projects such as the Millennium Run [108], Bolshoi [61], and the Outer Rim Simulation [55] have simulated the evolution of large-scale structures and have been widely used to generate realistic extragalactic foregrounds, including the Sunyaev-Zeldovich (SZ) effect, the Cosmic Infrared Background (CIB), and galaxy distributions [100, 109, 88].
Although N-body simulations can capture realistic correlations and non-Gaussian features, they are computationally expensive, and therefore only a limited number of microwave-sky realizations are currently available (e.g., [100, 109, 88]). Recently, [51] presented 500 high-resolution, full-sky millimeter-wave deep-learning (DL) simulations (“mmDL” simulations [^1] https://portal.nersc.gov/project/cmb/data/generic/mmDL/). These simulations are based on the “S10” simulation [100] and employ multiple neural-network architectures, including Generative Adversarial Networks (GANs [47]), to generate new realizations that preserve the statistical properties of the original “S10” data.
In this work, we use the extragalactic foreground realizations provided by ”mmDL” simulation. The simulation are provided at 30 GHz, 90 GHz, 148 GHz, 219 GHz, 277 GHz and 350GHz, we therefore interpolate the maps to the frequencies we need. Note that these foregrounds are only used for mocking LAT temperature observation since they are largely unpolarized under current sensitivity, except for the polarized point sources which have been proved to be sensitive to the PGW detection [94, 118, 119]. Besides, as described in [51], the flux of point sources at 148GHz has been truncated to 7 mJy, above which can be safely removed by point-source removal procedure under current and future sensitivity. Since the influence of polarized point sources is expected to be subdominant for the LAT configuration considered in this work [96], we leave a more detailed assessment to future work.
The input unlensed CMB maps are Gaussian realizations generated from a specific power spectrum obtained using the Boltzmann code CAMB [68], based on the Planck 2018 best-fit cosmological parameters [11] with a tensor-to-scalar ratio $r=0$. The lensing potential maps are converted from the convergence maps given by ”mmDL” simulation. We then utilize the Lenspyx [28, 93] package, which implements an algorithm to distort the primordial signal based on the lensing deflection ($d_{LM}=\sqrt{L(L+1)}\phi_{LM}$) generated above.
@@
where $N_{\text{white}}$ represents the white noise component, while $N_{\text{red}}$, $\ell_{\text{knee}}$, and $\alpha_{\text{knee}}$ characterize the contribution from $1/f$ noise. The total noise maps for LATs are generated as Gaussian realizations from the total noise power spectrum (with $\ell_{\text{max}}=6143$), using the parameters listed in Table 1. The beam-corrected noise power spectra of the LATs are shown in Fig. 3.
@@
[…]